# Characterization of a quantum bus between two driven qubits

### v1

@@ V Practical realization
Although the scheme described in this manuscript is rather generic, a scheme for realizing it in the context of bismuth or phosphorous donor based silicon spin qubits coupled to an optomechanical resonator is discussed in Ref. [28].
@@ VI Conclusions
In this work, we thoroughly analyze the use of driven qubits coupled to a harmonic oscillator to implement a $\sqrt{i\mathrm{SWAP}}$-gate. In many experimental realizations, the characteristic frequencies of the bare qubits and the bus differ by several orders of magnitude. In such cases, dressing the qubits via a driving field provides a mechanism for selectively coupling the qubits to the bus. First, we investigate a qubit readout mechanism based on the measurement of resonance frequency of the harmonic oscillator [31, 32, 16]. We show that the qubit state readout mechanism is more robust for the dressed qubit. Next, we have studied how the qubit-harmonic oscillator coupling can be leveraged to implement a two-qubit gate. We have analyzed the influence of various model parameters on the gate fidelity, identifying an optimal parameter regime for quantum computation.
@@ Acknowledgements.
This work was funded by the Finnish Quantum Flagship (project no. 210000621611) and the Research Council of Finland project DYNCOR (project no. 354735). We acknowledge grants of computer capacity from the Finnish Grid and Cloud Infrastructure (persistent identifier urn:nbn:fi:research-infras-2016072533). H. L. was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 852428).
Appendix A Resonance frequency shift
If the drive amplitude (Rabi frequency) $\Omega_{R}$ is close to resonance with the harmonic oscillator, we may apply the rotating wave approximation to the qubit-harmonic oscillator to simplify the coupling term. Working in the dressed qubit frame of reference, the Hamiltonian becomes:
$\hat{H}_{\mathrm{dress}}=\frac{\Omega_{R}}{2}\hat{\sigma}_{z}+\omega_{h}\hat{a}^{\dagger}\hat{a}+g(\hat{\sigma}_{-}\hat{a}^{\dagger}+\hat{\sigma}_{+}\hat{a})\;.$ (A.1)
The coupling term in (A.1) conserves the total number of quanta of the system, and it can be exactly diagonalized in the $\{|n,1\rangle,|n+1,0\rangle\}$ subspaces [52].
Hamiltonian (A.1) can be further simplified in the dispersive regime, where the detuning between the dressed qubit and the harmonic oscillator $\Delta_{R}=\Omega_{R}-\omega_{h}$ is large compared to the coupling strength $|\Delta_{R}|\gg g$ and $\langle\hat{a}^{\dagger}\hat{a}\rangle\ll\Delta_{R}^{2}/(4g^{2})$ [19, 53]. Applying a Schrieffer-Wolff transformation [54] of the form $\hat{U}=e^{g/\Delta_{R}(\hat{a}\hat{\sigma}_{+}-\hat{a}^{\dagger}\hat{\sigma}_{-})}$ provides the Hamiltonian in the new frame,
$\hat{H}^{\prime}_{\mathrm{dress}}=\left(\frac{\Omega_{R}}{2}+\frac{g^{2}}{2\Delta_{R}}\right)\hat{\sigma}_{z}+\left(\omega_{h}+\frac{g^{2}}{\Delta_{R}}\hat{\sigma}_{z}\right)\hat{a}^{\dagger}\hat{a}.$ (A.2)
In this regime, the qubit and the harmonic oscillator do not exchange energy, but the resonance frequency of the harmonic oscillator gets shifted by ±$\pm\frac{g^{2}}{\Delta_{R}}$ depending on the state of the qubit. This shift of the resonance frequency can be used as a readout mechanism to measure the state of the qubit [28, 31, 32, 16].
Appendix B Quantum bus between two dressed qubits
In this section, we consider two driven qubits coupled to a harmonic oscillator. The harmonic oscillator works as a quantum bus between qubits, entangling both qubits [19]. The Hamiltonian for the system can readily be obtained by generalizing Eq. (A.1),
$\hat{H}_{\mathrm{dress}}=\sum_{j}\frac{\Omega_{R,j}}{2}\hat{\sigma}_{z}^{j}+\omega_{h}\hat{a}^{\dagger}\hat{a}+\sum_{j}g_{j}(\hat{\sigma}_{-}^{j}\hat{a}^{\dagger}+\hat{\sigma}_{+}^{j}\hat{a})\;,$ (B.1)
where $j=1,2$ is the qubit index. Following the same steps as in Appendix A, applying a Schrieffer-Wolff transformation of the form $\hat{U}=e^{\sum_{j}g_{j}/\Delta_{R,j}(\hat{a}\hat{\sigma}_{+}^{j}-\hat{a}^{\dagger}\hat{\sigma}_{-}^{j})}$, in the dispersive regime the Hamiltonian acquires an effective qubit-qubit interaction term:
$\hat{H}^{\prime}_{\mathrm{dress}}=\sum_{j}\left(\frac{\Omega_{R,j}}{2}+\frac{g_{j}^{2}}{2\Delta_{R,j}}\right)$ $\hat{\sigma}_{z}^{j}+\left(\omega_{h}+\sum_{j}\frac{g_{j}^{2}}{\Delta_{R,j}}\hat{\sigma}_{z}^{j}\right)\hat{a}^{\dagger}\hat{a}$ (B.2)
$+\sum_{j}\frac{g_{j}g_{\bar{j}}}{2\Delta_{R,j}}(\hat{\sigma}_{+}^{j}\hat{\sigma}_{-}^{\bar{j}}+\hat{\sigma}_{-}^{j}\hat{\sigma}_{+}^{\bar{j}})\;,$
where $\bar{j}$ is the qubit opposite to $j$.
If the transition frequencies and the coupling constants of both qubits are the same, the interaction term (second line of Eq. (B.2)) commutes with the noninteracting Hamiltonian part $\hat{H}_{0}$, so the time evolution operator is given by [36, 55, 56]
[…]
If the qubit-qubit interaction time is fixed to $t=\frac{\pi|\Delta_{R}|}{4g^{2}}$ and the harmonic oscillator remains in the ground state, $\hat{U}(t)$ corresponds to a $\sqrt{\mp i\mathrm{SWAP}}$-gate depending on the sign of the detuning $\Delta_{R}$ [see Eq. (11)], up to 1-qubit gates. The interaction time can be controlled by tuning the drive amplitudes $\Omega_{R,j}(t)$ of the qubits, such that an appropriate pulse length tunes the dressed qubits in resonance with the harmonic oscillator. The $\sqrt{i\mathrm{SWAP}}$-gate is equivalent to a CNOT gate up to 1-qubit gates, so the interaction in Hamiltonian (B.2) is sufficient for universal quantum computation [5].

### v2

@@ V Practical realisation
Although the scheme described in this manuscript is rather generic, a scheme for realizing it in the context of bismuth or phosphorous donor-based silicon spin qubits coupled to an optomechanical resonator is discussed in Ref. [28].
@@ VI Conclusions
In this work, we thoroughly analyze the use of driven qubits coupled to a harmonic oscillator to implement a $\sqrt{i\mathrm{SWAP}}$-gate. In many experimental realisations, the characteristic frequencies of the bare qubits and the bus differ by several orders of magnitude. In such cases, dressing the qubits via a driving field provides a mechanism for selectively coupling the qubits to the bus. First, we investigate a qubit readout mechanism based on the measurement of the resonance frequency of the harmonic oscillator [33, 34, 16]. We show that the qubit state readout mechanism is more robust for the dressed qubit than the bare qubit as the steady state in the former generally exhibits information on both initial qubit states, and it is sufficient to fix the measurement time by observing the behaviour of the different amplitudes of the peaks. In contrast, for bare qubits, the resonance peak describing the steady state may mask the peak associated with the initial transient state. Moreover, dressed qubits tolerate thermal occupations better. We have also studied how the qubit-harmonic oscillator coupling can be leveraged to implement a two-qubit gate. We have analyzed the influence of various model parameters on the gate fidelity, identifying an optimal parameter regime for quantum computation.
@@
One key feature of using dressed qubits is the balanced steady state. A bare qubit is strongly affected by environmental damping, leading to relaxation into its ground state $|0\rangle$. In contrast, a dressed qubit relaxes into a mixed state, symmetrizing the $|0\rangle$ and $|1\rangle$ logical states. The state readout time may exceed the qubit’s decoherence time, so having a mixed steady state helps prevent a qubit flip due to thermal relaxation, thus improving the efficiency of initial state determination. However, most quantum algorithms require pure states, making the steady state of a dressed qubit unsuitable for qubit initialisation. A potential solution is to suppress the dressing amplitude $\Omega_{R}$ and apply a Zeeman field along the $z$-direction, enabling the qubit to decay into its bare ground state via thermal relaxation.
@@ Acknowledgements.
This work was funded by the Finnish Quantum Flagship (project no. 210000621611) and the Research Council of Finland (project no. 354735 and 321416). We acknowledge grants of computer capacity from the Finnish Grid and Cloud Infrastructure (persistent identifier urn:nbn:fi:research-infras-2016072533). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 852428).
Methods
Code availability
The underlying code for this study is available in GitLab and can be accessed via this link https://gitlab.jyu.fi/jyucmt/quantum-bus-between-driven-qubits.
Data availability
The datasets generated and analysed during the current study are available in the GitLab repository https://gitlab.jyu.fi/jyucmt/quantum-bus-between-driven-qubits.
Author contributions
AH developed the code, generated the results, and conducted the analysis presented in this work. All authors contributed to the discussion of the results and the writing of the manuscript.
Competing interests
All authors declare no financial or non-financial competing interests.